# Huge Ensembles Part II: Properties of a Huge Ensemble of Hindcasts Generated with Spherical Fourier Neural Operators

### v2

@@ Abstract
In Part I, we created an ensemble based on Spherical Fourier Neural Operators. As initial condition perturbations, we used bred vectors, and as model perturbations, we used multiple checkpoints trained independently from scratch. Based on diagnostics that assess the ensemble’s physical fidelity, our ensemble has comparable performance to operational weather forecasting systems. However, it requires several orders of magnitude fewer computational resources. Here in Part II, we generate a huge ensemble (HENS), with 7,424 members initialized each day of summer 2023. We enumerate the technical requirements for running huge ensembles at this scale. HENS precisely samples the tails of the forecast distribution and presents a detailed sampling of internal variability. For extreme climate statistics, HENS samples events 4 away from the ensemble mean. At each grid cell, HENS improves the skill of the most accurate ensemble member and enhances coverage of possible future trajectories. As a weather forecasting model, HENS issues extreme weather forecasts with better uncertainty quantification. It also reduces the probability of outlier events, in which the verification value lies outside the ensemble forecast distribution.
@@ 1 Introduction
Ensemble forecasts are an invaluable tool in weather and climate forecasting. By characterizing the probability distribution of possible future outcomes, they improve decision-making in the face of uncertainty (Mankin et al., 2020). In operational numerical weather prediction, there has been significant progress in accurately representing uncertainty to create probabilistic ensemble predictions (Palmer, 2002; Leutbecher and Palmer, 2008). In climate science, large ensembles of climate simulations are essential to separate internal variability from forced trends (Deser et al., 2020; Kay et al., 2015). Notable examples include the Community Earth System Model Large Ensemble with 100 members, and the 1000-member Observational Large Ensemble, which uses bootstrap resampling and signal processing methods to statistically characterize internal variability (McKinnon et al., 2017).
@@
Recent studies detecting and attributing extreme events state that small sample sizes are a major challenge to characterizing and researching extremes in the observational record (Thompson et al., 2017; Zhang et al., 2024; Bercos-Hickey et al., 2022; Philip et al., 2022). These low-likelihood high-impact events, such as heatwaves beyond 3 standard deviations from the climatological mean (e.g. those characterized in Zhang and Boos (2023)), have significant implications for human society and health. The observational record is limited to approximately fifty years, based on the start of the satellite era. Ensembles of weather and climate simulations alleviate this challenge by providing a large sample size of plausible atmospheric states and trajectories. The UNprecedented Simulated Extreme ENsemble (UNSEEN) approach tests whether an ensemble prediction system is fit for purpose by assessing its stability and fidelity (Kelder et al., 2022b, a). After they have been validated, ensembles have been used to quantify risk (Thompson et al., 2017), return-time (Leach et al., 2024), trends (Kirchmeier-Young and Zhang, 2020), attribution to climate change (Leach et al., 2021), and future changes (Swain et al., 2020) of low-likelihood, high impact extremes.
@@
Given the benefits of ensemble predictions, a core design decision is the number of ensemble members used in the ensemble. The ensemble size has significant implications for the forecast’s accuracy, uncertainty, and reliability. For climate simulations, Milinski et al. (2020) outline key steps and requirements to create initial condition ensemble. They present a framework to calculate the necessary number of ensemble members based on a user-specified maximum threshold for acceptable error and uncertainty. In weather forecasting, Leutbecher (2018) discuss the sampling uncertainty associated with finite sample sizes. They also assess how probabilistic scores, such as the continuous ranked probability score, converge as a function of ensemble size. Using reliability diagrams, Richardson (2001) demonstrate that 50-member ensembles are more reliable than 20-member ones. Siegert et al. (2019) assess the effect of ensemble size if the forecast is converted to a normal distribution, using the mean and standard deviation from the ensemble. In ensembles of up to 32 members, Buizza and Palmer (1998) characterize the effect of ensemble size on two metrics: (1) the skill of the best ensemble member at each gridpoint and (2) the outlier statistic. These two metrics reveal the benefit of larger ensembles, and we calculate these two statistics in a modern-day huge ensemble.
@@
Due to their computational costs, it is impractical to run massive ensembles with numerical weather and climate models. In Part I, we introduced an ensemble based on Spherical Fourier Neural Operators, Bred Vectors, and Multiple Checkpoints trained from scratch (SFNO-BVMC). The ensemble is orders of magnitude faster than comparable numerical simulations, so it enables the generation of huge ensembles (HENS). We validate the fidelity of SFNO-BVMC extremes using mean, spectral, and extreme diagnostics. Here, we use SFNO-BVMC to generate ensemble hindcasts initialized on each day of summer 2023. We choose summer 2023 as our test period because it is the hottest summer in the observed record (Esper et al., 2024). Therefore, it is an important period to validate forecasts of extreme heatwaves and to analyze low-likelihood high-impact heatwaves in a warming world. We use SFNO-BVMC to generate a huge ensemble with 7,424 members; each ensemble member is run for 15 days. Throughout this manuscript, we refer to this particular set of huge ensemble hindcasts during summer 2023 as "HENS." We contextualize the size of HENS compared to other large ensembles in Figure 1. HENS has more members than the majority of ensembles used in the past.
@@
We present specific properties of HENS as a tool to study extremes. To demonstrate the behavior across all of summer 2023, we calculate aggregate metrics, and we supplement a case study for extreme climate statistics and for weather forecasting. We organize the paper as follows:
@@
We list the technical requirements and considerations of creating ensemble sizes at this scale
@@
We illustrate how HENS enables improved climatic understanding of extreme weather events and their statistics
HENS provides a large sample size of counterfactual heat extremes
HENS produces better analogs to observed events as evidenced by at least 50% reduction in the best ensemble member RMSE metric and its sampling uncertainty
@@
We assess HENS’s ability to forecast extreme heat events
HENS finely samples the tails of the forecast distribution with many ensemble members
HENS reduces sampling uncertainty and results in a narrower confidence interval for extreme temperature forecasts
@@ 2.1 Technical Setup
In Part I, we used bred vectors to represent initial condition uncertainty, and we trained twenty-nine SFNO checkpoints from scratch to represent model uncertainty. Here, we combine these two perturbation techniques to create a 7,424-member ensemble: 256 initial condition perturbations for each of the 29 SFNO checkpoints. Bred vectors, which are flow-dependent perturbations, are generated separately for each SFNO checkpoint and each forecast initial time. Given the massive scale of the ensemble, we assess the ensemble spread created by bred vectors and multi-checkpointing. In Section F.2, we assess the exchangeability of SFNO checkpoints, and we show that at a lead time of 10 days, the checkpoints are exchangeable with one another.
@@
Generating and analyzing ensemble simulations at this scale is an important technical frontier. This quantity of data challenges the capabilities of existing climate model analysis workflows. The latitude and longitude dimensions are based on the 0.25-degree horizontal resolution of European Center for Medium-range Weather Forecasts Reanalysis v5 (ERA5), which is the SFNO-BVMC training dataset (Hersbach et al., 2020). The HENS simulation output is saved in 32-bit floating-point format. We store 12 variables and create a total dataset size of approximately 2 petabytes. Each variable in HENS thus takes 173 TB of space. Generating the ensemble for each initial date requires 21 TB of space and 45 minutes using 256 80GB NVIDIA A100 GPUs. With 92 initial times, generating HENS costs 18,432 GPU-hours (256 GPUs for 3 days). In Appendix B, we outline our post-processing strategy to analyze such large data volumes.
@@
To create the HENS hindcasts, we run SFNO in inference mode, in which Pytorch’s automatic differentiation is turned off. We use NVIDIA’s earth2mip library for inference. Our computations run on Perlmutter (Per, ), the high-performance computer at the National Energy Research Scientific Computing Center (NERSC). For optimal I/O, the ensemble is written to scratch on Perlmutter’s Lustre solid state scratch file system. At the time of generation, our scratch disk storage allocation was 100 TB, which is not large enough to hold all 2 petabytes of the ensemble at once. Scratch is meant to serve as a temporary high-bandwidth file storage solution, where data can be stored for up to 8 weeks. Therefore, the HENS simulations need to be transferred to larger, long-term storage on NERSC’s Community File System (CFS). We use Globus to transfer the ensemble from scratch to CFS. Since Globus leverages load-optimized, parallel file transfers via GridFTP (Ananthakrishnan et al., 2014), it is much faster than native Linux commands for copying data. The transfers run on dedicated nodes on Perlmutter and occur concurrently with ensemble generation.
Next, we present an example of the ensemble generation workflow. We first generate the ensemble initialized on June 1, 2023 at 00:00 UTC. After it completes, the Globus transfer from scratch to CFS begins. While this transfer is occurring, we begin generating the huge ensemble initialized on June 2, 2023 at 00:00 UTC. After the June 1 ensemble transfer completes, it is deleted from scratch. Crucially, the Globus transfers have 25 GB/s speeds, so they transfer each initial date’s 21 TB ensemble in approximately 15 minutes. With a native Linux command, the transfer took approximately 4 hours. Unlike the Linux commands, Globus transfers the ensemble in less time than it takes to generate the ensemble for the current initial date (45 minutes). Using simultaneous data generation and data transfer, we do not exceed our scratch disk storage allocation, since there is never more than two initial days’ worth of ensembles stored at once. Our computational allocation goes entirely towards data generation rather than data transfer, and GPUs are never idle while the transfer occurs.
@@ 2.2 Regenerating the Ensemble
ML’s computational efficiency can lead to a different paradigm of data transfer. Instead of sharing the simulation output, it may be more practical to share the trained ML model weights.111ML model weights refer to the learned parameters of the architecture: during training, an ML model uses an optimization algorithm to update these weights to minimize the loss function. At the end of training, the ML model weights can be shared and readily used for inference. This removes the requirement to save the data on a public, high-bandwidth storage location. By sharing the model weights, other scientific users can regenerate the entire ensemble or a particular subset of members that simulate a particular extreme event of interest.
To ensure reproducibility, we designed HENS based on a random seed. During inference, this random seed determines the initial condition perturbation, which is created by initially adding spherical random noise to Z500 to start the breeding cycle (see Part I). During inference, we set the seed for the random number generator. After the completion of the HENS simulation, we validate that we can regenerate specific ensemble members. We open-source our ML model weights and inference code for others to regenerate ensemble members of interest. We save the 12 variables listed in Table 1. One could regenerate specific ensemble members to save more variables to assess other atmospheric phenomena during the simulation. In this way, ML model weights provide a data compression mechanism to avoid storing the data produced from massive ensembles. However, there are two shortcomings of the seed-based reproducibility. First, the random-number-generation approach may change in the future as Pytorch and other scientific packages are upgraded. Second, non-determinism on individual GPUs (e.g. bit-level soft errors, in which the in-memory state of the model could change during the simulation) or in specific frameworks (e.g. Vonich and Hakim (2024))) pose a challenge for reproducibility. Although we can reproduce our simulations on our computing environment, bit-for-bit reproducibility is an important direction for future research, especially across different hardware and computing systems. To assess whether either of these two problems has occurred, checksums of the original ensemble output can be used to ensure that the ensemble regeneration occurred successfully.
@@ 3.1 Sampling the Forecast Distribution with Huge Ensembles
In Part I, we validated a 58-member SFNO-BVMC ensemble using the continuous ranked probability score (CRPS), spread-error ratio, threshold-weighted CRPS (twCRPS), reliability diagram, and Receiver Operating Characteristic (ROC) curve. On these metrics, SFNO-BVMC performs comparably to the Integrated Forecast System (IFS), a leading ensemble numerical weather prediction model based on traditional physics solvers. These metrics show that SFNO-BVMC has a realistic, reliable, and well-calibrated ensemble spread. These are prerequisite characteristics for running huge ensembles. If they are not met, the forecast distribution would not provide useful probabilistic information. Sampling the tails of such a forecast distribution would not yield realistic estimates of internal variability. In this section, we analyze the sampling properties of the HENS forecasts.
We explore the large sample behavior of HENS at a 10-day lead time. We present three reasons for this choice of lead time. First, at 10 days, the spread-error ratio has reached approximately 1 for all variables (shown in Figure 8 of Part I), indicating that the ensemble has a reasonable representation of its uncertainty. Second, 10-day forecasts can be validated against observations because they are within the predictability limit of approximately 14 days. The 10-day ensemble mean root mean squared error (RMSE) is still lower than the climatology RMSE. Therefore, the HENS trajectories are still dependent on the initial conditions, and they are realistic possible outcomes from the initial conditions. After the predictability limit is breached, the ensemble distribution should more closely represent the climatological distribution. In this scenario, the model becomes free-running, since its forecast is no longer directly tied to the initial conditions. At 10 days, we can still compare the forecasts to the observations, and this comparison serves as a powerful way to validate the predictions from new entirely ML-based weather models. Third, after 10 days, the ensemble trajectories have diverged due to uncertainty in synoptic-scale atmospheric motion. This greater ensemble dispersion allows for a more thorough characterization of different future outcomes from the initial time.
@@
If the tails of the forecast distribution were Gaussian, sampling 4 standard deviations events from the ensemble mean would require approximately 7,000 members. Section F.3 theoretically derives the expected information gain for samples from a Gaussian distribution. The theoretical Gaussian gain is shown by the dotted line in Figure 2. We use this estimate of 7,000 to guide our choice of ensemble size for HENS. We are constrained by available memory, computational resources, and data movement time from scratch to rotating disk file storage. Based on these constraints, we determine that we have the computational budget for 7,424 members: 256 members each for 29 trained SFNO checkpoints. (See Section 2.2 of Part I for more information. an empirical justification of our choice of 29 checkpoints.)
@@
Then . We consider the following variables: 2m temperature, total column water vapor, 850hPa temperature, 500hPa geopotential, and 10m zonal wind. Due to the nature of the instantaneous atmospheric flow, each ensemble member predicts mean conditions at some locations and extreme conditions at other locations. At a given time, there will likely be extreme conditions occurring somewhere on Earth, simply due to the spatial variation of weather. In our calculation of information gain, we do not consider the spatial distribution of extremes, which varies significantly within each ensemble member. Instead, we wish to assess the distribution across ensemble members. Therefore, we calculate the information gain on the global land mean values of each ensemble member. This allows us to assess the ensemble members in aggregate and how far each ensemble member is from the ensemble mean.
@@
Figure 2 shows the information gain as a function of ensemble size. The 7,424-member ensemble has an information gain of 4. This means that HENS is large enough to have at least one ensemble member that is 4 standard deviations away from the ensemble mean (on average). For all variables, the HENS gain closely follows the theoretical Gaussian gain. This result is not completely surprising: averaging over a large number of grid cells implies that a Central Limit Theorem should apply (even though the grid box values are neither independent nor identically distributed), wherein the global land averages behave similarly to a Gaussian random variable. We use this theoretical Gaussian behavior to inform our choice of ensemble size. In any case, with such a large information gain for the full ensemble, we can be confident that HENS is an effective way to sample the tails of the forecast distribution, and it includes ensemble members that simulate trajectories of low-likelihood events. To sample ensemble members that are 5 standard deviations away from the mean, ensemble members are necessary according to the Gaussian approximation. This ensemble size would require a large compute expenditure and inline ensemble pruning. Since it would not be practical to save the entire ensemble, it would be necessary to only save the ensemble members that simulate the rarest of extreme events (Webber et al., 2019). We also visualize the information gain at each grid cell in Figure 3. At the grid-cell level, there is a deviation from the Gaussian behavior. In the HENS ensemble, some grid cells depart from the expected Gaussian value of approximately 4. Similarly, in a 50-member ensemble, the grid cells depart from the expected Gaussian value of approximately 2. At all locations, HENS is better able to sample extreme events, compared to a 50-member ensemble.
@@
Using the full HENS distribution, we calculate the ensemble mean, standard deviation, and 0.1, 10, 90, and 99.9 percentiles of global land-mean 2m temperature. (Note that the 0.1 and 99.9 percentiles represent the 1000-day extreme low and high thresholds, respectively.) Figure 4 compares the degree to which smaller ensemble sizes can accurately emulate these statistics. For each ensemble size, 2,000 bootstrap random samples are taken from HENS, and the resulting statistics and their uncertainty are shown relative to the corresponding statistic from the full ensemble. We provide a detailed description of calculating these statistics in Section F.4.3. In Section F.4.1 and Figure 19, we present Gaussian theory for how the percentiles should change with ensemble size . We use this theory to calculate the "analytic uncertainty" dotted lines in Figure 4. In Section F.4.2, we use extreme value theory to calculate theoretical estimates of the extreme percentiles in Figure 4. Additionally, in Section F.5, we discuss how the normalized uncertainty is calculated. Because of the seasonal cycle in the ensemble spread across summer 2023 (not shown), we normalize the statistics by the uncertainty from the full HENS forecast at each initial date. This enables comparison of the sampling characteristics across all forecasts from summer 2023.

### v3

@@ Abstract
In Part I, we created an ensemble based on Spherical Fourier Neural Operators. As initial condition perturbations, we used bred vectors, and as model perturbations, we used multiple checkpoints trained independently from scratch. Based on diagnostics that assess the ensemble’s physical fidelity, our ensemble has comparable performance to operational weather forecasting systems. However, it requires orders of magnitude fewer computational resources. Here in Part II, we generate a huge ensemble (HENS), with 7,424 members initialized each day of summer 2023. We enumerate the technical requirements for running huge ensembles at this scale. HENS precisely samples the tails of the forecast distribution and presents a detailed sampling of internal variability. HENS has two primary applications: (1) as a large dataset with which to study the statistics and drivers of extreme weather and (2) as a weather forecasting system. For extreme climate statistics, HENS samples events 4 away from the ensemble mean. At each grid cell, HENS increases the skill of the most accurate ensemble member and enhances coverage of possible future trajectories. As a weather forecasting model, HENS issues extreme weather forecasts with better uncertainty quantification. It also reduces the probability of outlier events, in which the verification value lies outside the ensemble forecast distribution.
@@ 1 Introduction
Ensemble forecasts are an invaluable tool in weather and climate forecasting. By characterizing the probability distribution of possible future outcomes, they improve decision-making in the face of uncertainty (Mankin et al., 2020). In operational numerical weather prediction, there has been significant progress in accurately representing uncertainty to create probabilistic ensemble predictions (Palmer, 2002; Leutbecher and Palmer, 2008). In climate science, large ensembles are essential to separate internal variability from forced trends (Deser et al., 2020; Kay et al., 2015). Notable examples include the 100-member Community Earth System Model Large Ensemble and the 1000-member Observational Large Ensemble, which uses bootstrap resampling and signal processing methods to represent internal variability (McKinnon et al., 2017).
@@
Small sample sizes are a major challenge to characterizing and researching extremes in the observational record (Thompson et al., 2017; Zhang et al., 2024; Bercos-Hickey et al., 2022; Philip et al., 2022). Low-likelihood high-impact events, such as heatwaves beyond 3 standard deviations away from the climatological mean (e.g. those characterized in Zhang and Boos (2023)), have significant implications for human society and health. The observational record is limited to approximately fifty years, based on the start of the satellite era. Ensembles of weather and climate simulations alleviate this challenge by providing a large sample size of plausible atmospheric states and trajectories. Finkel et al. (2023) characterize sudden stratospheric warming events using a large dataset of subseasonal-to-seasonal ensemble hindcasts. The UNprecedented Simulated Extreme ENsemble (UNSEEN) approach tests whether an ensemble prediction system is fit for purpose by assessing its stability and fidelity (Kelder et al., 2022b, a). After they have been validated, ensembles have been used to quantify risk (Thompson et al., 2017), return-time (Leach et al., 2024), trends (Kirchmeier-Young and Zhang, 2020), attribution to climate change (Leach et al., 2021), and future changes (Swain et al., 2020) of low-likelihood, high-impact extremes.
@@
Given the benefits of ensemble predictions, a core design decision is the number of members used in the ensemble. The ensemble size has significant implications for the forecast’s accuracy, uncertainty, and reliability. For climate simulations, Milinski et al. (2020) outline key steps and requirements to create initial condition ensembles. Based on a user-specified threshold for acceptable error and uncertainty, they present a framework to calculate the required number of members. In weather forecasting, Leutbecher (2018) discuss the sampling uncertainty associated with finite sample sizes. They also assess how probabilistic scores, such as the continuous ranked probability score (CRPS), converge as a function of ensemble size. Using reliability diagrams, Richardson (2001) demonstrate that 50-member ensembles are more reliable than 20-member ones. Siegert et al. (2019) assess the effect of ensemble size if the forecast is converted to a normal distribution, using the mean and standard deviation from the ensemble. In ensembles of up to 32 members, Buizza and Palmer (1998) characterize the effect of ensemble size on two metrics: (1) the skill of the best ensemble member at each grid point and (2) the outlier statistic. These two metrics reveal the benefit of larger ensembles, and we calculate these two statistics in a modern-day huge ensemble.
@@
Due to computational costs, it is impractical to run massive ensembles with numerical weather and climate models. In Part I (Mahesh et al., 2024), we introduced an ensemble based on Spherical Fourier Neural Operators, Bred Vectors, and Multiple Checkpoints trained from scratch (SFNO-BVMC). The ensemble is orders of magnitude faster than comparable numerical simulations, so it enables the generation of huge ensembles (HENS). We validated the fidelity of SFNO-BVMC extremes using mean, spectral, and extreme diagnostics. Here in Part II, we use SFNO-BVMC to generate ensemble hindcasts initialized on each day of summer 2023. We choose summer 2023 as our test period because it is the hottest summer in the observed record (Esper et al., 2024). Therefore, it is an important period to validate forecasts of extreme heatwaves and to analyze low-likelihood high-impact heatwaves in a warming world. We use SFNO-BVMC to generate a huge ensemble with 7,424 members; each ensemble member is run for 15 days. Throughout this manuscript, we refer to this particular set of huge ensemble hindcasts during summer 2023 as "HENS." We contextualize the size of HENS compared to other large ensembles in Figure 1. HENS has more members than the majority of ensembles used in the past.
@@
We present specific properties of HENS as a tool to study extremes. To demonstrate the behavior across all of summer 2023, we calculate aggregate metrics, and we provide a case study for extreme climate statistics and for weather forecasting. In this manuscript, we present three contributions:
@@
We list the technical requirements and considerations of creating ensemble sizes at this scale.
@@
As a large hindcast of simulated weather extremes, HENS enables improved climatic understanding of extreme weather events and their statistics.
HENS samples low-likelihood events at the tail of the forecast distribution.
HENS provides a large sample size of counterfactual heat extremes.
HENS produces better analogs to observed events by reducing the error of the best ensemble member
@@
We assess HENS’s ability to forecast extreme heat events.
HENS provides a finer sampling of the conditional forecast distribution, given that the forecast values exceed a climatological extreme threshold
HENS reduces sampling uncertainty and has narrower confidence intervals for extreme temperature forecasts
@@ 2.1 Technical Setup
In Part I, we used bred vectors to represent initial condition uncertainty, and we trained twenty-nine SFNO checkpoints from scratch to represent model uncertainty. Here, we combine these two perturbation techniques to create a 7,424-member ensemble: 256 initial condition perturbations for each of the 29 SFNO checkpoints. Bred vectors, which are flow-dependent perturbations, are generated separately for each SFNO checkpoint and each forecast initial time.
@@
Generating and analyzing ensemble simulations at this scale is an important technical frontier. The latitude and longitude dimensions are based on the 0.25-degree horizontal resolution of European Center for Medium-range Weather Forecasts Reanalysis v5 (ERA5) (Hersbach et al., 2020), which is the SFNO-BVMC training dataset. The HENS simulation output is saved in 32-bit floating-point format. We store 12 variables and create a total dataset size of approximately 2 petabytes. Each variable in HENS thus takes 173 TB of space. Generating the ensemble for each initial date requires 21 TB of space and 45 minutes using 256 80GB NVIDIA A100 GPUs. With 92 initial times, generating HENS costs 18,432 GPU-hours (256 GPUs for 3 days). This quantity of data challenges the capabilities of existing climate model analysis workflows. In Appendix B, we outline our post-processing strategy to analyze such large data volumes. If we stored all the channels in the ensemble (not our chosen subset of 12 channels), then the total size of the dataset would be approximately 25 petabytes.
@@
To create the HENS hindcasts, we run SFNO in inference mode, in which Pytorch’s automatic differentiation is turned off. We use NVIDIA’s earth2mip library for inference. Our computations run on Perlmutter (Per, ), the high-performance computer at the National Energy Research Scientific Computing Center (NERSC). For optimal I/O, the ensemble is written to scratch on Perlmutter’s Lustre solid state scratch file system. At the time of generation, our scratch disk storage allocation was 100 TB, which is not large enough to hold all 2 petabytes of the ensemble at once. Scratch is meant to serve as a temporary high-bandwidth file storage solution, where data can be stored for up to 8 weeks. Therefore, the HENS simulations need to be transferred to larger, long-term storage on NERSC’s Community File System (CFS). We use Globus to transfer the ensemble from scratch to CFS. Since Globus leverages load-optimized, parallel file transfers via GridFTP (Ananthakrishnan et al., 2014), it is much faster than native Linux commands for copying data. The transfers run on dedicated nodes on Perlmutter and occur concurrently with ensemble generation. Due to the computational efficiency of ML weather forecasts, it is feasible to generate 256 ensemble members (each running on 1 GPU) in parallel per minute. This introduces new data transfer considerations to ensure the data can be moved to its storage location in time.
We present an example of the ensemble generation workflow. We first generate the ensemble initialized on June 1, 2023 at 00:00 UTC. After it completes, the Globus transfer from scratch to CFS begins. While this transfer is occurring, we begin generating the huge ensemble initialized on June 2, 2023 at 00:00 UTC. After the June 1 ensemble transfer completes, it is deleted from scratch. Crucially, the Globus transfers have 25 GB/s speeds, so they transfer each initial date’s 21 TB ensemble in approximately 15 minutes. With a native Linux command, the transfer took approximately 4 hours. Unlike the Linux commands, Globus transfers the ensemble in less time than it takes to generate the ensemble (45 minutes). Using simultaneous data generation and data transfer, we do not exceed our scratch disk storage allocation. At any given time, no more than two initial days’ worth of ensembles are stored on scratch. Our computational allocation goes entirely towards data generation rather than data transfer, and GPUs are not idle while the transfer from scratch to CFS occurs.
@@ 2.2 Regenerating the Ensemble
ML’s computational efficiency can lead to a different paradigm of data transfer. Instead of sharing the simulation output, it may be more practical to share the trained ML model weights111ML model weights refer to the learned parameters of the architecture: during training, an ML model uses an optimization algorithm to update these weights to minimize the loss function. At the end of training, the ML model weights can be shared and readily used for inference. and the initial conditions. This removes the requirement to save the data on a public, high-bandwidth storage location. By sharing the model weights, other scientific users can regenerate the entire ensemble or a particular subset of members that simulate a particular extreme event of interest.
To ensure reproducibility, we designed HENS based on a random seed. During inference, this random seed determines the initial condition perturbation, which is created by initially adding spherical random noise to Z500 to start the breeding cycle (see Part I). During inference, we set the seed for the random number generator. After the completion of the HENS simulation, we validate that we can regenerate specific ensemble members. We open-source our ML model weights and inference code for others to regenerate ensemble members of interest. We save the 12 variables listed in Table 1. One could regenerate specific ensemble members to save more variables to assess other atmospheric phenomena during the simulation. In this way, ML model weights provide a data compression mechanism to avoid storing the data produced from massive ensembles. However, there are two shortcomings of the seed-based reproducibility. First, the random-number-generation approach may change in the future as Pytorch and other scientific packages are upgraded. Second, non-determinism on individual GPUs (e.g. bit-level soft errors, in which the in-memory state of the model could change during the simulation) or in specific frameworks (Vonich and Hakim, 2024) pose a challenge for reproducibility. Although we can reproduce our simulations on our computing environment, bit-for-bit reproducibility is an important direction for future research, especially across different hardware and computing systems. To assess whether either of these two problems has occurred, checksums of the original ensemble output can be used to ensure that the ensemble regeneration occurred successfully.
@@ 3.1 Sampling the Forecast Distribution with Huge Ensembles
In Part I, we validated a 58-member SFNO-BVMC ensemble using the CRPS, spread-error ratio, threshold-weighted CRPS (twCRPS), reliability diagram, and Receiver Operating Characteristic (ROC) curve. On these metrics, SFNO-BVMC performs comparably to the Integrated Forecast System (IFS), a leading ensemble numerical weather prediction model based on traditional physics solvers. These metrics show that SFNO-BVMC has a realistic, reliable, and well-calibrated ensemble spread. These are prerequisite characteristics for running huge ensembles. If they are not met, the forecast distribution would not provide useful probabilistic information. Sampling the tails of such a forecast distribution would not yield realistic estimates of internal variability. In this section, we analyze the sampling properties of the HENS forecasts.
We explore the large sample behavior of HENS at a 10-day lead time. We present three reasons for this choice of lead time. First, at 10 days, the spread-error ratio has reached approximately 1 for all variables (shown in Figure 8 of Part I), indicating that the ensemble has a reasonable representation of its uncertainty. Second, 10-day forecasts can be validated against observations because they are within the predictability limit of approximately 14 days. The 10-day ensemble mean root mean squared error (RMSE) is still lower than the climatology RMSE. Therefore, the HENS trajectories are still dependent on the initial conditions, and they are realistic possible outcomes from the initial conditions. After the predictability limit is breached, the ensemble distribution should more closely represent the climatological distribution. In this scenario, the model becomes free-running, since its forecast is no longer directly tied to the initial conditions. At 10 days, we can still directly compare the forecasts to the observations, and we do not have to rely on comparison of the climatological characteristics of the free-running model (e.g. comparison of the probability density functions across space and time). Third, after 10 days, the ensemble trajectories have diverged due to uncertainty in synoptic-scale atmospheric motion. This greater ensemble dispersion allows for a more thorough characterization of different future outcomes from the initial time.
@@
If the tails of the forecast distribution were Gaussian, sampling 4 standard deviation events from the ensemble mean would require approximately 7,000 members. Section G.3 theoretically derives the expected information gain for samples from a Gaussian distribution. The theoretical Gaussian gain is shown by the dotted line in Figure 2. We use this estimate of 7,000 to guide our choice of ensemble size for HENS. We are constrained by available memory, computational resources, and data movement time from scratch to rotating disk file storage. Based on these constraints, we determine that we have the computational budget for 7,424 members: 256 members each for 29 trained SFNO checkpoints. (See Section 2.2 of Part I.)
@@
Then . We consider the following variables: 2m temperature, total column water vapor, 850hPa temperature, 500hPa geopotential, and 10m zonal wind. Due to the nature of the instantaneous atmospheric flow, each ensemble member predicts mean conditions at some locations and extreme conditions at other locations. At a given time, there will likely be extreme conditions occurring somewhere on Earth, simply due to the spatial variation of weather. We do not consider the spatial distribution of extremes, which varies significantly within each ensemble member. Instead, we wish to assess the distribution across ensemble members. Therefore, we calculate the information gain on the global land mean values of each ensemble member. This allows us to assess the ensemble members in aggregate and how far each ensemble member is from the ensemble mean.
@@
Figure 2 shows the information gain as a function of ensemble size. The 7,424-member ensemble has an information gain of 4. This means that HENS is large enough to have at least one ensemble member that is 4 standard deviations away from the ensemble mean (on average). For all global land-means of these variables, the HENS gain closely follows the theoretical Gaussian gain. This result is not completely surprising: averaging over many grid cells implies that a Central Limit Theorem should apply (even though the grid box values are neither independent nor identically distributed), wherein the global land averages behave similarly to a Gaussian random variable. We use this theoretical Gaussian behavior to inform our choice of ensemble size. With such a large empirical information gain, HENS includes ensemble members that simulate trajectories of low-likelihood events. To sample ensemble members that are 5 standard deviations away from the mean, ensemble members would be necessary according to the Gaussian approximation. This ensemble size would require a large compute expenditure and inline ensemble pruning. Since it would not be practical to save the entire ensemble, it would be necessary to only save the ensemble members that simulate the rarest of extreme events (Webber et al., 2019).
To complement our analysis of the gain of the global land means, we also assess the gain of huge ensembles at each grid cell. While the gain of the global-land mean value closely follows the Gaussian approximation, there are significant deviations from Gaussianity at the local level. In Figure 3, we visualize the information gain at each grid cell. In the HENS ensemble, some grid cells depart from the expected Gaussian value of approximately 4. Similarly, in a 50-member ensemble, the grid cells depart from the expected Gaussian value of approximately 2.
@@
Using the full HENS distribution, we calculate the ensemble mean, standard deviation, and 0.1, 10, 90, and 99.9 percentiles of global land-mean 2m temperature. (Note that the 0.1 and 99.9 percentiles represent the 1000-day extreme low and high thresholds, respectively.) Figure 4 assesses the ability of different ensemble sizes to emulate these statistics accurately. For each ensemble size, we take 2,000 bootstrap random samples from HENS, and the resulting statistics and their uncertainty are shown relative to the corresponding statistic from the full ensemble. We provide a detailed description of calculating these statistics in Section G.4.3. In Section G.4.1 and Figure 24, we present Gaussian theory for how the percentiles should change with ensemble size . We use this theory to calculate the "analytic uncertainty" dotted lines in Figure 4. In Section G.4.2, we use extreme value theory to calculate theoretical estimates of the extreme percentiles in Figure 4. Additionally, in Section G.5, we discuss how the normalized uncertainty is calculated. Because of the seasonal cycle in the ensemble spread across summer 2023, we normalize the statistics by the uncertainty from the full HENS forecast at each initial date. This enables comparison of the sampling characteristics across all forecasts from summer 2023.